### v1

# Accreditation Against Limited Adversarial Noise

Abstract
I present an easily-applicable method for upgrading any pre-existing digital accreditation (a variety of quantum verification) protocol - that assumes IID CPTP error - to function correctly when error is modelled as being adversarial - albeit slightly modified to more closely resemble physical reality - with no diminution in their efficiency or suitability for near-term usage.
I Introduction
Accreditation has provided an efficient, scalable, and practical method for quantifying the the quality of particular computations. This is a vital requirement for making use of quantum computers in the NISQ era [1], when quantum computations will be unreliable due to interactions with the surrounding environment - known as noise [2] - that induce erroneous - i.e. unintended / not in the circuit being executed - operators in a computation; so some notion of whether the outputs of quantum computations are usable - for whatever purpose they are required for - is needed.
I note that accreditation is not the only method of evaluating the quality of quantum computations and/or quantum computers. One such class of protocols is randomized benchmarking [3, 4, 5, 6]. These protocols do not aim to measure the how well a quantum computer can implement a specific computation - e.g. preparing a six-qubit Greenberger-Horne-Zeilinger (GHZ) state [7] - but rather how well it performs benchmarking computations - which usually comprise of a series of gates that, if implemented with no error, return a known measurement outcome with certainty - and hence obtain a measure of the quality of the quantum computer itself, instead of a measure of the quality of a computation’s outputs like accreditation does.
A more useful measure, in the long term, is how well specific computations are implemented, and hence to what extent data obtained from quantum devices can be trusted. Verification [8, 9, 10, 11] is a large field with many protocols aimed at solving exactly this problem. These protocols typically run a quantum computation, allowing for the possibility of error, and - via some machinations - decide whether to accept or reject the returned outputs. This is a more limited measure of the quality of a specific computation than the ideal-actual variation distance (as defined in Def. 1) - which accreditation protocols provide.
Definition 1.
For any circuit, $\mathcal{C}$, and an execution of it, $\tilde{\mathcal{C}}$, the ideal-actual variation distance of that execution (denoted as $\nu\big{[}\tilde{\mathcal{C}}\big{]}$) is the variation distance between the probability distribution the execution would sample from if there were no error (i.e. the ideal case) - which I denote by $\mathcal{O}_{I}\big{[}\mathcal{C}\big{]}$ and the distribution it actually samples from.
Furthermore, verification protocols, broadly speaking, rely on a different kind of assumption to accreditation. Verification has typically used the cryptographic framework of delegated computation  [^1] This is also known as an interactive prover system, wherein successful verification protocols have been shown to exist for a very large set of problems [12, 13], although - ironically - noise may severely restrict when interactive prover systems exist [14].  [15] (i.e. is based on Alice and Bob interacting and precluding error typically [^2] There exists verification protocols that do not have this assumption, that use two quantum devices [16, 17, 18], but here I focus on single-device protocols. either in state preparation [19, 20, 21] or measurement [22, 23, 17, 24]); accreditation has instead tended to base its assumptions on the physics within quantum computers. By turning from the cryptographic assumptions of the verification protocols [8], from which it originates, to physics-based assumptions; accreditation has been able to develop assumptions, based on experimental evidence, and then protocols [25], using the aforementioned assumptions, to initiate a vein of research that has, thus far: provided the aforementioned scalable protocol to achieve confidence in computation outputs, proven itself experimentally implementable and loosened its assumptions [26], and directly led to the first method for quantifying the correctness of the outputs of quantum analogue simulation [27].
But these successes have come at a cost. Most relevantly, for the purposes of this paper, accreditation has hitherto required abandoning the adversarial noise model present in many verification protocols [8]; moving, instead, to a model where noise induces identically and independently distributed (IID) CPTP error [^3] Error which may be represented by a completely positive trace preserving map acting on both the system and its environment.. Returning to an adversarial error model has proven difficult as the physics-based assumptions clash with the adversarial model of noise. This is resolved herein with a “hybrid” approach, using a limited form of adversarial noise wherein the cryptographic setting / Alice and Bob formalism [15], as used in adversarial noise models, is used but modified with a protocol (Protocol 1) for how quantum circuits are executed that both Alice and Bob participate in - still leaving Bob to be as malicious as he likes but limiting, based on the aforementioned physics-based assumptions, his knowledge and abilities. This permits me to improve upon the IID CPTP error model, while retaining validity in all situations where the IID assumption was valid. The result is protocols with less stringent assumptions, loosening the requirements on the physical devices used to implement them.
This is achieved through exploiting limits placed on Bob (based on experimental realities - that the probability of error varies little in executions of the same circuit, on the same hardware, in the same environment - and inspired by single-qubit gates experiencing gate-independent error in pre-existing accreditation protocols [25, 26]) allowing methods to overcome the factors that prohibited adversarial noise in extant accreditation protocols.
This paper proceeds, from here, with an introduction to the error model used - in Sec. II.1 - and its justification - in Sec. II.2. I then present my main result: the adaptation of accreditation protocols to the newly-developed error model in Sec. III. The paper then concludes - in Sec. IV - with a discussion of further possible developments.
II Adversarial Error Model
II.1 Error Model Presentation
My adversarial error model is an adaption of the cryptographic setting [25]. In the cryptographic setting, Alice is attempting to get Bob to perform a sampBQP [28, 29, 30] quantum computation [^4] Meaning a computation to obtain a sample from a specified distribution, that can be efficiently performed on a quantum computer. Note that sampBQP includes both BQP and fBQP (the set of functions efficiently computable on a quantum computer). and provide her with the true results, while only having the ability to perform polynomially-bounded classical computation herself. Bob is computationally unbounded so, if he wished, could implement the computation Alice requires but his aim is instead to trick Alice by providing her with the results of an incorrect computation and to have her believe he has acted honestly i.e. that the results she receives are the outputs of the computation she wanted to be performed when they are not.
My specific error model is slightly modified from the traditional error model used in the cryptographic setting, mainly through limitations on Bob. These limitations are entirely contained in how the circuits Alice requests be executed are executed, which is presented in Protocol 1, but this first requires Def. 2.
Definition 2.
A gate in a circuit is said to be redacted if the gate’s position in the circuit - i.e. when it is applied and to which qubits - is specified but the operator it represents - e.g. if it is a Pauli $\hat{X}$ gate or Pauli $\hat{Z}$ gate - is not.
A final concept required to be formally defined before the presentation of how the circuits Alice requests be executed are executed is a notion of the unique choice of CPTP maps representing error in an execution.
Definition 3.
A CPTP set corresponding to a specific circuit is a possible set of the CPTP maps that represent the error occurring in the execution of that circuit.
I can then present the protocol by which sets of circuits, under my error model, are executed, in Protocol 1.
  1. 1.
    Alice provides the set of circuits (differing only in their single-qubit gates), $\mathcal{S}$, to be executed.
  2. 2.
    Bob receives the full details of $\mathcal{S}$, with all single-qubit gates redacted, and picks a set of CPTP sets (each element of which is a valid CPTP set for any element of $\mathcal{S}$) - denoted $\Xi$ subject to the condition that, for all CPTP sets in $\Xi$, the resulting probability of error in the execution of any circuit in $\mathcal{S}$, with any element of $\Xi$ representing its error, is within the interval:
    $\big{[}P_{0}\big{(}1-\beta\big{)},P_{0}\big{(}1+\beta\big{)}\big{]},$ (1)
    for a known (to Alice and Bob, in advance) $\beta\in\mathbb{R}^{+}$.
  3. 3.
    For each circuit in $\mathcal{S}$:
    1. (a)
      Bob chooses a CPTP set, $\xi$, from $\Xi$.
    2. (b)
      The circuit is executed with $\xi$ determining the error occurring during the execution.
    3. (c)
      Alice and Bob are both given the measurement outcomes of the execution.
Protocol 1 Formal Adversarial Model: How Sets of Circuits are Executed
Alice is free to initiate Protocol 1 whenever and as often as she likes; Bob is free to act how he likes to achieve his stated malicious aims but must execute Protocol 1 as prescribed - but is free to act nefariously within the permitted bounds - when it is invoked.
II.2 Physical Origins of the Error Model
The two main limitations on the adversarial error / Bob are:
  1. 1.
    The single-qubit gates are redacted when the circuits are shown to Bob (in step 2 of Protocol 1).
  2. 2.
    For any set of circuit executions within a single use of Protocol 1, the respective probabilities of error occurring are all contained in the interval $[P_{0}(1-\beta),P_{0}(1+\beta)]$ (where $\beta\in\mathbb{R}^{+}$ is known to Alice, but $P_{0}$ is not) i.e. that the probability of error is only slightly different in different executions of similar circuits in quick succession. This is represented in Protocol 1 by Eqn. 1.
The first limitation corresponds to single-qubit gates experiencing gate-independent error. This is a standard assumption the the pre-existing accreditation protocols (in Refs. [25, 26, 27]) and follows from single-qubit gates typically being the least error-prone components of a quantum computer [31, 32] (which has held true over time [33, 34], and can be most clearly seen in Ref [35, Fig.5]): the error is so small that the error in different single-qubit gates does not differ much. This is a “standard [assumption] in the literature on noise characterisation and mitigation” [36] and has seen extensive use in theoretical work [3, 37, 38, 39, 40, 41, 42, 43, 44, 45, 36].
The second limitation is a weakening of the IID assumption in previous accreditation protocols. The intuition behind it is that the same hardware executing very similar circuits in quick succession - as happens in Protocol 1 will experience similar error in each circuit as:
  1. 1.
    The external factors affecting error have to be fluctuating over a very specific time-span - slow enough to not fluctuate much within the execution of circuits but fast enough to change between executions of a circuit - to cause significant differences in the error in different executions executed in quick succession.
  2. 2.
    The hardware executing the similar circuits is the same in each execution and any error-inducing aspects of the hardware are unlikely to change much in the short time between executions.
  3. 3.
    The effect of any aspect of the circuits being executed that may change the error is minimized, as the circuits are very similar.
Although no paper has, to my knowledge, sought to directly validate this limitation, it can also be justified experimentally:
  1. 1.
    Ref. [26, Figure. 5] ran the accreditation protocol - developed therein - many times, producing a probability of error in the traps for each use of the protocol. The set of error probabilities generated by this can be seen to vary little across many uses of the protocol.
  2. 2.
    Ref. [46, Fig. 6, Fig. 7, and Fig. 8] investigated the error rates in single-qubit and two-qubit gates. It found these error rates are very rarely far from their average. Note that the variation shown in this paper is over a much longer time-span (days) than this limitation requires (seconds).
  3. 3.
    Ref. [35, Fig. 9] examined NISQ devices and - across many devices - plotted the error rates of different qubits in each device. Fig. 9 shows the error bars on the error rate to be - with some notable exceptions - small fractions of the error rate and varying much more across qubits than for a fixed qubit over time.
  4. 4.
    Ref. [47, Fig. 1(b)] looked at the error rates of two-qubit gates on differing pairs of qubits, it shows that - again, over a time-span of days - for a specific pair of qubits, the days when the error rate deviates far from its average are rare, with more extreme deviations being rarer.
Finally, this limitation can be seen as an aim of quantum computer hardware engineering: as quantum computers improve and their actual outputs approach the errorless outputs, with decreasing variance, the typical difference - by any measure - between two executions of similar circuits will tend to zero.
To my knowledge, I am the first to explicitly state and use this second limitation. However, I note that it is already implicitly accepted in the community by the acceptance of randomized benchmarking [3, 4, 5, 6] as a meaningful measure: if even the same circuit repeated multiple times, in very quick succession, produces wildly varying probabilities of error then the variation of error probabilities implies that previous measures mean almost nothing for future computations or the quality of a quantum device over a meaningful time-span. This is not the case, and Ref. [47, Fig. 6] shows that, over a span of days, the results of randomized benchmarking do not vary over a very large range.
There is, however, another, subtle, restriction on Bob: he controls the CPTP maps acting on all executions, but does not control any stochastic elements of the implementation of the CPTP map. I.e. he may specify that a specific CPTP map in the execution applies stochastic Pauli error, but he does not decide if this results in a Pauli $\hat{X}$ gate being applied. Furthermore, Bob is not informed of the outcomes of these stochastic processes - in spite of choosing which stochastic processes occur. This represents that the environment - where any memory of past error that the noise represented by Bob has must be - is ironically not, consistently or over long time-spans, protected from noise acting on it and destroying any information/memory stored in it, so any hostile intelligence that lurks in the noise must have a comparatively short and small memory  [^5] Quantum hardware engineers have expended much effort to keep the the information inside quantum computers intact, so there is no reason to assume the environment can - with non-negligible probability - store significant amounts of information without extensive engineering effort to enable it to do so.. This argument is strengthened by the potential use of error suppression techniques such as dynamic decoupling [48].
III Accreditation Protocol
III.1 Trap and Target Circuits
In this paper, I do not propose to develop new trap circuits or target circuits. In fact, I would prefer to make minimal changes to pre-existing protocols. To this end, I will talk in generalities about traps and targets, and only specify conditions on target circuits and trap circuits - met by all target and trap circuits in the accreditation protocols in Ref. [25], Ref. [26], and Ref. [27] - that they must meet. I will assume that I have two efficient classical algorithms, $P_{\textit{targ}}$ and $P_{\textit{trap}}$, that, if given a valid quantum circuit as input, return the corresponding target and a matching trap, respectively.
  1. 1.
    In target circuits and trap circuits all error occurring is twirled (via Pauli twirls or some other twirl) to stochastic error.
  2. 2.
    Each trap acts as a Bernoulli trial with the success probability being the probability error occurs in that trap according to the particular CPTP maps - representing the error - in the trap.
  3. 3.
    If no error occurs in a trap, it gives a specific output, $m$, and if error does occur the trap does not give the output $m$ with probability at least $k\in[0,1]$.
  4. 4.
    Target circuits and trap circuits differ only in their single-qubit gates.
The above assumptions have mentioned twirling CPTP error to stochastic error, which I define formally in Def. 4. This is an important step in all accreditation protocols, as it reduces the error to a known, more easily quantified form.
Definition 4.
CPTP error within a quantum circuit is said to be twirled to stochastic error if, via the addition of only single-qubit gates to the circuit, the error is effectively transformed to stochastic error, without otherwise affecting the outputs of the circuit. Likewise, gates are said to be twirled if any error occurring in them is twirled. An important example is Pauli twirling [49, 50, 51], where the resulting stochastic error consists exclusively of Pauli operators.
My new adversarial error model may, however, require a slight modification of the trap circuits and target circuits: the error model they are designed for may not consider “hiding” the outputs as there was no adversary to “hide” them from. With my new adversarial error model, this becomes necessary as otherwise Bob may be able to identify traps and targets. The required hiding / encrypting of outputs is achieved via Algorithm 2.
$\mathbb{Input:}$
$\bullet$ A circuit, $\mathcal{C}$, to generate traps or a target for.
$\bullet$ Two algorithms, $P_{\textit{trap}}$ and $P_{\textit{targ}}$, as detailed above.
$\bullet$ A Boolean, labelled isTarget, denoting if a trap or target is to be generated.
  1. 1.
    If isTarget == True:
    1. (a)
      $\mathcal{C}^{\prime}$ = $P_{\textit{targ}}\big{(}\mathcal{C}\big{)}$.
    Else:
    1. (a)
      $\mathcal{C}^{\prime}$ = $P_{\textit{trap}}\big{(}\mathcal{C}\big{)}$.
  2. 2.
    Generate a random bit string with the same length as the number of measurements in $\mathcal{C}^{\prime}$, referred to as the key.
  3. 3.
    For measurement, M, in $\mathcal{C}^{\prime}$:
    1. (a)
      Calculate the single-qubit unitary, $\mathcal{U}_{M}$, that, if applied immediately before measurement, M, would flip the outcome.
    2. (b)
      If ( the bit in the key corresponding to measurement M ) == 1:
      1. i.
        Add $\mathcal{U}_{M}$ to $\mathcal{C}^{\prime}$ immediately before measurement M.
$\mathbb{Return}:$ $\mathcal{C}^{\prime}$ and the key.
Algorithm 2 Generating Traps and Targets with Hidden Outputs
I note that the required single-qubit unitaries, $\mathcal{U}_{M}$, will exist for any single-qubit measurement and that the outcome of the circuit - if the single-qubit gates of the circuit are not known, as is the case for Bob as they are redacted for him - is irretrievable from the measurement outcomes without the key. This hiding of the circuit outputs is information-theoretically secure (i.e. has perfect security) [52], assuming the single-qubit gates are redacted, and comparable to the one-time pads in universal blind quantum computing [53]. Henceforth, I will denote $P_{\textit{targ}}$ and $P_{\textit{targ}}$, with the changes presented above to make their outputs hidden, by $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$ respectively.
A similar method may be used to hide the inputs into circuits.
III.2 Presentation of the Upgraded Accreditation Protocol
With trap and target circuits established, I can consider the full accreditation protocol to be developed. In line with Sec. III.1, for Sec. III.2 I assume I already have some method of generating trap and target circuits conforming to the requirements in Sec. III.1. However, before the full protocol can be presented, in Protocol 3, I must establish the statistical foundations of the new protocol - in Sec. III.2.1 - and the core mechanics of the protocol - in Sec. III.2.2.
III.2.1 Statistical Basis of the New Accreditation Protocol
Before presenting the accreditation protocol, it is first useful to present a purely statistical lemma (Lemma 2) that will later enable the accreditation protocol. This presentation first begins with Def. 5 and Lemma 1.
Definition 5.
For any set of real values , =$\mathcal{R}=\big{\{}r_{j}\big{\}}^{|\mathcal{R}|}_{j=1}$, define: $\underline{\mathrm{Avg}\big{(}\mathcal{R}\big{)}}=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}r_{j}\bigg{)}.$
Lemma 1.
Given a set of real values, $\mathcal{R}\subset[(1-\beta)P_{0},(1+\beta)P_{0}]$; if $\beta\in\mathbb{R}$ is known, $P_{0}$ can be bounded as:
$\dfrac{\mathrm{Avg}(\mathcal{R})}{1+\beta}$ $\leq P_{0}\leq\dfrac{\mathrm{Avg}(\mathcal{R})}{1-\beta}.$ (2)
Proof.
Denote =$\mathcal{R}=\{r_{j}|1\leq j\leq|\mathcal{R}|\}$, then:
$\mathrm{Avg}(\mathcal{R})$ $=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}r_{j}\bigg{)}.$ (3)
As $\forall r_{j}\in\mathcal{R}$, $r_{j}\in[(1-\beta)P_{0},(1+\beta)P_{0}]$; $\forall r_{j}\in\mathcal{R}$, $(1-\beta)P_{0}\leq r_{j}\leq(1+\beta)P_{0}$. Therefore,
$(1-\beta)P_{0}$ $=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}(1-\beta)P_{0}\bigg{)}$ (4)
$\leq\mathrm{Avg}(\mathcal{R})$ (5)
$\leq\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}(1+\beta)P_{0}\bigg{)}=(1+\beta)P_{0}$ (6)
$\Rightarrow\dfrac{\mathrm{Avg}(\mathcal{R})}{1+\beta}$ $\leq P_{0}\leq\dfrac{\mathrm{Avg}(\mathcal{R})}{1-\beta}.$ (7)
I may then present and prove Lemma 2, which makes use of Lemma 1. Whereas Lemma 1 was concerned with characterizing the interval $\mathcal{R}$, Lemma 2 aims to place a bound on any value occurring within $\mathcal{R}$.
Lemma 2.
Given a set of real values, $\mathcal{R}\subset[(1-\beta)P_{0},(1+\beta)P_{0}]$; if $\beta\in[0,1)$ is known, $\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$,
$y\leq\bigg{(}1+2\beta\bigg{)}\mathrm{Avg}(\mathcal{R}).$ (8)
Proof.
Using Lemma 1,
$P_{0}\leq\dfrac{\mathrm{Avg}(\mathcal{R})}{1-\beta}.$ (9)
$\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$, $y\leq(1+\beta)P_{0}$. Therefore, $\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$,
$y\leq(1+\beta)P_{0}\leq\dfrac{1+\beta}{1-\beta}\mathrm{Avg}(\mathcal{R}).$ (10)
Using the Taylor series of $\dfrac{1}{1-\beta}$, which holds for all values $\beta$ can take, and neglecting quadratic (or higher) terms:
$\dfrac{1+\beta}{1-\beta}\approx 1+2\beta\Rightarrow y\leq\bigg{(}1+2\beta\bigg{)}\mathrm{Avg}(\mathcal{R}),$ (11)
where I have assumed that the approximation is tight enough to not affect the inequality. ∎
III.2.2 Core Mechanics of the New Accreditation Protocol
Before its formal presentation, in Sec. III.2.3, the core mechanics of the accreditation protocol are presented. This takes the form of Lemma 3 and Lemma 4.
Lemma 3.
If given a classical polynomial-time algorithm, $P^{\prime}_{\textit{trap}}$, generating traps corresponding to a circuit, $\mathcal{C}$, and the corresponding $m$, and $k$ (as defined in Sec. III.1); assuming the probability of error in each execution of a trap is independent of the outcomes of all preceding traps, that any trap detects any specific error with probability at least $k\in[0,1]$, and all traps are executed via Protocol 1, $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps are sufficient to estimate, $\mathrm{Avg}(\mathcal{R})$ (where $\mathcal{R}$ is the set of probabilities of error in each trap), to within error $\theta$ with confidence $\alpha$.
Proof.
Due to the independence, between traps, of whether error occurs, as assumed, $\mathrm{Avg}(\mathcal{R})$ can be estimated, as required, using Hoeffding’s inequality [54] and $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps, as the fraction of traps returning an incorrect measurement outcome serves as an approximation to the probability of the traps not measuring the correct results, to required error. This then provides the required bound - to required error - by dividing it by $k$. ∎
With Lemma 3 established, I can present and prove a central theorem of the the accreditation protocol’s core mechanics: Lemma 4.
Lemma 4.
If given a polynomial algorithm, $P_{\textit{trap}}^{\prime}$, generating traps for a fixed circuit, $\mathcal{C}$, and the corresponding $m$, and $k$ (as defined in Sec. III.1); assuming the probability of error in each execution is independent of the outcomes of all preceding executions, that any trap detects any specific error with probability at least $k\in[0,1]$, and all traps are executed via Protocol 1, the probability of error in any execution of any circuit executed within the same singular use of Protocol 1 - such as a target circuit  [^6] As the probability of error in the target would, due to the assumptions made, be within $[(1-\beta)P_{0},(1+\beta)P_{0}]$. - as the traps can be upper bounded by:
$\dfrac{1+2\beta}{k}\big{(}\bar{v}+\theta\big{)},$ (12)
using $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps, where: $\bar{v}$ is the fraction of traps giving an incorrect measurement outcome, $\theta\in\mathbb{R}^{+}$ may be chosen arbitrarily, $\beta$ is as in Protocol 1, and $\alpha$ is the confidence required in the bound.
Proof.
As in Lemma 3, define the set $\mathcal{R}$ to be the probability of error occurring in each specific trap in a set of traps executed, along with any number of other circuits (that are assumed to meet the conditions to be executed alongside the traps in Protocol 1), in a singular use of Protocol 1. This implies that the probability of error in any specific trap mentioned above - and in any circuit executed within the same single use of Protocol 1 - is within the interval:
$[P_{0}(1-\beta),P_{0}(1+\beta)].$ (13)
Therefore, due to Lemma 2, for any circuit executed in the aforementioned single use of Protocol 1, labelled circuit $j$, the probability of error occurring in its execution, $p_{j}$, is bounded by:
$p_{j}\leq\bigg{(}1+2\beta\bigg{)}\mathrm{Avg}(\mathcal{R}).$ (14)
The probability a specific trap returns an incorrect measurement outcome, given error occurs, is, by assumption of having the required traps, lower bounded by $k\in[0,1]$; hence, an upper bound on $\mathrm{Avg}(\mathcal{R})$ can be approximated, as in Lemma 3, to within error $\theta$, with confidence $\alpha$. Let $\bar{\nu}$ denote the approximation achieved from sampling using the traps (i.e. $\bar{\nu}$ is the fraction of traps measuring in incorrect output). Then, with confidence, $\alpha$,
$p_{j}\leq\dfrac{1+2\beta}{k}\big{(}\bar{v}+\theta\big{)}.$ (15)
As per Lemma 3, which is applicable due to the assumption that the probability of error is independent of previous measurement outcomes - this requires only $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps. ∎
III.2.3 Formal Protocol For Accreditation
The core components of my accreditation protocol have now been constructed and presented - in Sec. III.2.1 and Sec. III.2.2 - so I can now present my accreditation protocol formally, in Algorithm 3. This algorithm is presented as would be used by Alice, presupposing the error model in Sec. II.2.
$\mathbb{Input:}$
$\bullet$ A circuit, $\mathcal{C}$, meeting any required conditions.
$\bullet$ A required accuracy of the bound to output, $\theta$.
$\bullet$ A required confidence in the above accuracy, $\alpha$.
$\bullet$ The relevant $P^{\prime}_{\textit{targ}}$, $P^{\prime}_{\textit{trap}}$, $m$, and $k$.
$\bullet$ The value of $\beta$.
  1. 1.
    Calculate $N_{l}=\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}+1$.
  2. 2.
    $\{\mathcal{C}_{j}\}_{j=1}^{N_{l}}$ = $N_{l}$ circuits generated using $P_{\textit{trap}}$ on input $\mathcal{C}$.
  3. 3.
    $\mathcal{C}^{\prime}$ = circuit generated using $P_{\textit{targ}}$ on input $\mathcal{C}$.
  4. 4.
    Execute every circuit in $\{\mathcal{C}_{j}\}_{j=1}^{N_{l}}\cup\{\mathcal{C}^{\prime}\}$ in a random order using Protocol 1.
  5. 5.
    $\mathrm{TargResult}$ = output of executing $\mathcal{C}^{\prime}$
  6. 6.
    $\mathrm{TrapResult}$= fraction of $\{\tilde{\mathcal{C}}_{j}\}_{j=1}^{N_{l}}$ that does not output $m$.
$\mathbb{Return}:$ $\mathrm{TargResult}$, $\frac{1+2\beta}{k}\big{(}\mathrm{TrapResult}+\theta\big{)}$.
Algorithm 3 Formal Accreditation Protocol
The correct functioning of Algorithm 3, as an accreditation protocol, is proven in Theorem 1.
Theorem 1.
Assuming all sets of circuits executed are executed via Protocol 1, and that the classical algorithms, $P^{\prime}_{\textit{targ}}$ and $P^{\prime}_{\textit{trap}}$, generate trap and target circuits conforming to the conditions in Sec. III.1, then Algorithm 3 provides a sample from a probability distribution and a bound on the ideal-actual variation distance between that distribution and the distribution generated by the measurements in $\mathcal{C}$ when no error occurs (to within accuracy $\theta$, with confidence $\alpha$).
Proof.
All circuits executed in Algorithm 3 are executed within a single use of Protocol 1. Therefore, if each probability of error in a given trap is independent, then Lemma 4 implies that, as the target circuit is a circuit executed alongside $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps in a single use of Protocol 1, the probability of error in the target circuit’s execution, $p_{t}$, is upper bounded as:
$p_{t}\leq\bigg{(}\dfrac{1+2\beta}{k}\bigg{)}\big{(}\bar{v}+\theta\big{)},$ (16)
where $\bar{v}$ is the fraction of traps returning an incorrect measurement result and $k\in[0,1]$ is as defined in Sec. III.1. As mentioned, for Eqn. 16 to hold, it is required that the probability of error in each trap is independent of the outcomes of every prior trap. This is not initially guaranteed as Bob is free to choose probabilities based on previous measurement outcomes. However, due to the changes made to each trap and target in Algorithm 2, to hide the outputs of each trap from Bob with perfect security, Bob does not know the result of any trap or target. Therefore, he cannot make any of his choices based on the results of any prior trap or target  [^7] I additionally note that Bob cannot remove this independence by basing his choice of CPTP maps on error in previous traps as although Bob knows which CPTP maps are applied to all previous circuits executed, he does not know which stochastic errors have occurred due to them, only the probabilities of each occurring (due to having chosen the CPTP maps) because he doesn’t control any aspect of the CPTP maps’ applications, due to the construction of Protocol 1. Therefore, Bob cannot base the choice of CPTP maps in a trap on the outcome of any stochastic property of previous traps as they are unknown to him.. Hence, each trap can be considered to be independent i.e. the probability of error in each trap or target is the result of Bob’s choices but must remain completely independent of the results of any prior trap or target.
As per Protocol 1, when Bob is given the set of circuits to execute - which in this case would be the traps with the target hidden among them - the single-qubit gates are redacted: therefore, as the targets and the traps only differ in their single-qubit gates (as assumed in Sec. III.1) Bob cannot distinguish traps from targets at any stage in Protocol 3. The redaction additionally allows for the unbreakable encryption of the output of each trap or target, as in Algorithm 2, so Bob cannot locate the target via the outputs either. ∎
IV Discussion
In this paper I have presented an easily-applicable method for upgrading any pre-existing digital accreditation protocol to considering all error to be adversarial. This has necessitated the development of a new model of adversarial error, where Bob is limited according to experimentally extracted rules. These rules - that single-qubit gates are hidden from Bob and that the probability of error in similar circuits executed in quick succession have similar probabilities of error - allow any extant accreditation protocol to function almost identically when experiencing adversarial error as they do when the error is IID. Therefore, the desirable qualities of these protocols, such as not requiring excessively long circuits, or ancilla qubits, or extensive connectivity between qubits are preserved; leading to no diminution in their suitability for near-term usage.
The principle question left open is to what extent can the limits on Bob be relaxed. This would have to entail changes to Protocol 1. I believe the redaction of single-qubit gates cannot be eliminated without substantial changes to the trap, as otherwise Bob could distinguish target and trap circuits - although it is not clear trap and target circuits that are indistinguishable to Bob can be constructed without requiring another limitation on Bob. A slight improvement can be obtained using Ref. [26, Theorem 1], which allows the error in single-qubit gates to be weakly gate dependent: this could be added to my cryptographic model by allowing Bob to define error in the - still redacted - single-qubit gates that depends on the single-qubit gates (single-qubit gates would still have to be redacted to prevent Bob identifying the target) but the differing CPTP maps - for each single-qubit gate - must differ, in the diamond norm, by, at most, some small, known, value. This is a further step towards reality and the restriction on the differences between the gate-dependent error reflects the very small errors in single-qubit gates [35, Fig.5]. Finally I note that, if I take the philosophical position that error that does not affect the state of the qubits effectively does not occur and can be considered to have never been applied, I can relax the assumption that Bob does not directly control stochastic processes in the CPTP maps modelling error. This is as the Pauli twirling, accompanied by the random Hadamard gates, make it impossible for Bob to know whether gates he has applied have any affect - which, as mentioned above, can be interpreted as him not being able to decide if he has effectively applied the error. This breaks any ability of Bob to apply error in a way that is dependent on the error in previous circuit executions. Due to this, the error can be considered to be independent between circuit executions.
V Acknowledgements
I would like to thank Animesh Datta and Theodoros Kapourniotis for useful conversations.
This work was supported, in part, by a EPSRC IAA grant (G.PXAD.0702.EXP) and the UKRI ExCALIBUR project QEVEC (EP/W00772X/2).