### v2

# Accreditation Against Limited Adversarial Noise

Abstract
I present an accreditation (a variety of quantum verification) protocol where error is assumed to be adversarial (in contrast to the IID assumption used in previous accreditation protocols) - albeit slightly modified to more closely resemble physical reality - with no diminution in efficiency or suitability for near-term usage.
I Introduction
Accreditation provides an efficient and scalable method for quantifying the the quality of quantum computations, without trusting any aspect of the computations. This is a vital requirement for using quantum computers in the NISQ era [1], when quantum computations will be unreliable due to interactions with the surrounding environment - known as noise [2] - that induce erroneous operators in a computation; so some measure of whether the outputs of quantum computations are usable is needed.
I note that accreditation is not the only method of evaluating quantum computations or quantum computers. One such class of protocols is randomized benchmarking [3, 4, 5, 6]. These protocols do not aim to measure the how well a quantum computer can implement a specific computation - e.g. preparing a six-qubit Greenberger-Horne-Zeilinger (GHZ) state [7] - but rather how well it performs benchmarking computations - which usually comprise of a series of gates that, if implemented with no error, return a known measurement outcome with certainty - and hence obtain a measure of the quality of the quantum computer itself, instead of a measure of the quality of a computation’s outputs like accreditation does.
A more useful measure, in the long term, is how well specific computations are implemented and if data obtained from quantum devices can be trusted. Verification [8, 9, 10, 11] is a large field with many protocols aimed at solving exactly this problem. These protocols typically run a quantum computation, allowing for the possibility of error, and - via some machinations - decide whether to accept or reject the returned outputs. This is a more limited measure of the quality of a specific computation than the ideal-actual variation distance (as in Def. 1) which accreditation protocols provide.
Definition 1.
For any circuit, $\mathcal{C}$, and an execution of it, $\tilde{\mathcal{C}}$, the ideal-actual variation distance of that execution (denoted as $\nu\big{[}\tilde{\mathcal{C}}\big{]}$) is the variation distance between the probability distribution the execution would sample from if there were no error (i.e. the ideal case) - which I denote by $\mathcal{O}_{I}\big{[}\mathcal{C}\big{]}$ and the distribution it actually samples from.
Furthermore, verification protocols, broadly speaking, rely on a different kind of assumption to accreditation. Verification has typically used the cryptographic framework of delegated computation [12] (i.e. is based on Alice and Bob interacting and precluding error typically [^1] There exists verification protocols that do not assume there is no error in some particular aspect of a computation, using two quantum devices [13, 14, 15], but here I focus on single-device protocols. either in state preparation [16, 17, 18] or measurement [19, 20, 14, 21]); accreditation has instead tended to base its assumptions on the physics within quantum computers. By turning from the cryptographic assumptions of verification protocols [8], from which it originates, to physics-based assumptions; accreditation has been able to develop assumptions, and then protocols [22] - using those assumptions - to initiate a vein of research that has, thus far: provided the aforementioned scalable protocol to achieve confidence in computation outputs, proven itself experimentally implementable [23], and led to the first method for quantifying the quality of the outputs of quantum analogue simulation [24].
But these successes have come at a cost. Most relevantly, for this paper, accreditation has hitherto required abandoning the adversarial noise model present in many verification protocols [8]; moving, instead, to a model where noise induces identically and independently distributed (IID) CPTP error [^2] Meaning in each execution of a circuit, the error is represented by identical and independent - across multiple executions - completely positive trace preserving maps acting on both the system and its environment.. Returning to an adversarial error model / problem setting has proven difficult as the physics-based assumptions of accreditation protocols clash with the adversarial model of noise. This is resolved herein by upgrading the protocol in Ref. [23] to use/assume a limited form of adversarial noise wherein the cryptographic setting / Alice and Bob formalism [12], as used in adversarial noise models, is used but modified with a protocol (Protocol 1) for how quantum circuits are executed that both Alice and Bob participate in - still leaving Bob to be as malicious as he likes but limiting, based on the aforementioned physics-based assumptions, his knowledge and abilities. This permits me to improve upon the IID CPTP error model, in Ref. [23], while retaining validity in all situations where it was valid. The result is a protocol with less stringent assumptions, loosening the requirements on the physical devices used to implement them.
[…]
Error modelled as CPTP maps Error modelled as CPTP maps
Error is IID Error is adversarial (but limited)
GI single-qubit gate error GI single-qubit gate error
Table 1: Table showcasing the upgrade presented in this paper as a comparison between the assumptions in Ref. [23] and the protocol herein. Note that GI is shorthand for Gate-Independent (see Sec. II.4).
This is achieved through exploiting limits placed on Bob (based on experimental realities - that the probability of error varies little in executions of the same circuit, on the same hardware, in the same environment - and inspired by single-qubit gates experiencing gate-independent (GI) error in pre-existing accreditation protocols [22, 23]) allowing methods to overcome the factors that prohibited adversarial noise in extant accreditation protocols.
This paper proceeds, from here, with an introduction to the problem setting - in Sec. II.3 - and its justification - in Sec. II.4. I then present my main result: the adaptation of accreditation protocols to the newly-developed problem setting / model of error in Sec. III. The paper then concludes - in Sec. IV - with a discussion of further possible developments.
II Problem Setting of This Paper
As this manuscript bridges the gap between the physics-based assumption model of previous accreditation papers and the cryptographic setting that is more typical of verification protocols, I take an approach that I hope will satisfy both tribes: below I present the problem setting - where Alice, Bob, and their respective objectives are introduced - the crux of which is an protocol (Protocol 1) for how circuits are executed (which Alice and Bob both participate in), and then the physics of errors occurring in a circuit execution are used to support and justify the problem setting, for those who prefer a physics-based error model.
The foundation of my problem setting [^3] Which can be seen as akin to the error model in previous accreditation protocols. is the assumption that error in any operation (e.g. state preparation, gate implementation, or measurement) may be considered as the ideal/errorless operation followed or preceded by a CPTP map, as in Ref. [23]. This is justified as any map from and to density matrices is a CPTP map. Hence, any erroneous operation must be a CPTP map and so any erroneous map may be written as the errorless/ideal operation either followed or preceded by a CPTP map - as unitaries are CPTP maps.
II.1 Introduction to Alice and Bob
With its foundations established, I now present my problem setting, which is an adaption of the cryptographic setting [22]. In the cryptographic setting, Alice is attempting to get Bob to perform a sampBQP [25, 26, 27] quantum computation [^4] Meaning a computation to obtain a sample from a specified distribution, that can be efficiently performed on a quantum computer. Note that sampBQP includes both BQP and fBQP (the set of functions efficiently computable on a quantum computer). and provide her with the true results, while only having the ability to perform polynomially-bounded classical computation herself. Bob is computationally unbounded so, if he wished, could implement the computation Alice requires but his aim is instead to trick Alice by providing her with the results of an incorrect computation and to have her believe he has acted honestly i.e. that the results she receives are the outputs of the computation she wanted to be performed when they are not.
II.2 Problem Setting: Preliminaries and Definitions
My specific problem setting is slightly modified from the traditional cryptographic setting, mainly through limitations on Bob. These limitations are entirely contained in how the circuits Alice requests be executed - and Bob tries to corrupt - are executed, which is presented in Protocol 1 in Sec. II.3, but Protocol 1 first requires the concept of redaction (introduced in Def. 2) and a related concept (in Def. 3).
Definition 2.
A gate in a circuit is said to be redacted if the gate’s position in the circuit - i.e. when it is applied and to which qubits - is specified but the operator it represents is not. For an example of a circuit with a redacted gate, see Fig. 1.
Figure 1: An example circuit with a redacted gate (the block in the middle of the circuit). Note how the location of the gate - in terms of time and which qubits are affected - is depicted but which operation the gate represents is hidden.
Definition 3.
If given a circuit with all single-qubit gates redacted, the set of all circuits - with no redactions - that the given circuit (with redactions) could be (if the redactions were removed) is called its redaction class. For an example of circuits in the same redaction class, see Fig. 2.
Figure 2: Example circuits within the same redaction class, where the redaction class corresponds to the redacted circuit in Fig. 1. Note that the circuit in Fig. 1 could be any of these if the redaction were removed.
The second and final round of terminology required for the presentation of Protocol 1 is related to how I specify the CPTP maps that are modelling the error in a single execution of a circuit. This is performed via CPTP lists and they are defined in Def. 4.
Definition 4.
A CPTP list is an ordered set of CPTP maps.
However, a CPTP list is just a list of CPTP maps. In order to to be able to describe / determine the error in a given circuit, the CPTP list must contain exactly the right number of CPTP maps. In this case the CPTP list is said to fit the circuit.
Definition 5.
A CPTP list fits a circuit if it may be used to describe error  [^5] Any resulting error - including no error - is valid. It need not be in any way physically justified or correspond to a particular quantum computer’s typical error. in an execution of that circuit.
A CPTP list achieves this by containing exactly one CPTP map corresponding to each location in the circuit where error may occur (These locations are: immediately after gates, immediately after state prep, and immediately before measurement)  [^6] The mapping between the CPTP maps in the CPTP list and the locations error occurs in the circuit is made using the ordering of the CPTP list and matching this with an ordering applied to the locations in the circuit where error may occur..
A CPTP list will fit many different circuits and for any execution of a circuit there exists a CPTP list that both fits the circuit and accurately represents the error occurring in the circuit. I refer to this as the CPTP list determining the error occurring in a circuit execution.
For an example of a CPTP list determining the error in a circuit execution, see Fig. 3.
Figure 3: If $\Xi_{1}=[\xi_{1},\xi_{2},\xi_{3}]$ is a CPTP list used to determine the error in an execution of the top circuit in Fig. 2, then Fig. 3 depicts the circuit that is actually executed. $\xi_{1}$ describes the error due to state preparation, $\xi_{2}$ describes the error due to the single-qubit gate, and $\xi_{3}$ describes the error due to measurement.
Note that notation is slightly abused to display CPTP maps as gates (depicting error) in a circuit, to remedy this slightly they are highlighted in red to denote they are not unitaries but are CPTP maps.
Note 1.
For any set of circuits within the same redaction class, if a given CPTP list fits one of them, it fits all of them.
An important feature to note about CPTP lists and the circuits it fits is mentioned - without proof - in Lemma 1. Proof is omitted as Lemma 1 follows from the above discussion.
Lemma 1.
The error in an execution of a circuit is entirely and exclusively determined by the CPTP list used to describe its error (and the outcomes of any stochastic processes that may occur in the application of the error described by the CPTP list  [^7] As a CPTP map may describe the stochastic application of operations.).
Definition 6.
The probability of error in the execution of a circuit, due to a CPTP list, is the probability that the execution does not provide a sample from the same distribution that executing the circuit as intended (i.e. with no error applied) would. Due to Lemma 1, the probability of error is entirely determined by the CPTP list used to represent the error in the execution. For example, using $\Xi_{1}=[\xi_{1},\xi_{2},\xi_{3}]$ from Fig. 3, if, for all density matrices, $\rho$:
$\xi_{1}(\rho)$ $=\dfrac{1}{2}\rho+\dfrac{1}{2}\hat{X}\rho\hat{X}$ (1)
$\xi_{2}(\rho)$ $=\dfrac{1-0.01}{2}\rho+\dfrac{1+0.01}{2}\hat{X}\rho\hat{X}$ (2)
$\xi_{3}(\rho)$ $=\dfrac{1+0.01}{2}\rho+\dfrac{1-0.01}{2}\hat{X}\rho\hat{X},$ (3)
then the probability of error in the execution of a circuit where these CPTP maps (i.e. $\Xi_{1}$) determine the error is the probability that any of $\xi_{1}$, $\xi_{2}$, or $\xi_{3}$ cause error. In this case, the probability of error is $\approx 0.875$. This quantity is not always meaningful if the error in an execution is not stochastic Pauli error, therefore, for all the traps and target circuits used throughout this paper, the probability of error is defined to be the probability error occurs once the error has been effectively reduced to - and can be considered as - stochastic Pauli error (by the Pauli twirling, defined in Def. 8, and present in all traps and targets considered herein).
Definition 7.
A Set of Probabilistically Similar CPTP Lists with parameter $\beta\in\mathbb{R}$ (henceforth denoted as a PSCLβ) is a CPTP list such that for any CPTP list in the set, the probability of error in any circuit execution it determines the error for is within the interval:
$\big{[}P_{0}\big{(}1-\beta\big{)},P_{0}\big{(}1+\beta\big{)}\big{]},$ (4)
for some $P_{0}\in[0,1]$. Additionally, all CPTP maps in the PSCLβ must be of the same size, so if one fits a given circuit all CPTP maps in the PSCLβ do.
For example consider two CPTP lists: $\Xi_{2}=[\xi_{1}]$ and $\Xi_{3}=[\xi_{2}]$ (where $\xi_{1}$ and $\xi_{2}$ are as in Def. 6). Then the set $\{\Xi_{2},\Xi_{3}\}$ can be considered as a PSCLβ for any $\beta>0.01$ (where the value of $P_{0}$ in this PSCLβ is $0.5$).
With both PSCLβ and redaction classes defined, they may be related via the below Lemma 2. As will be seen in Protocol 1, Alice chooses a redaction class (implicitly, by choosing a subset of one to execute) and Bob chooses a PSCLβ to determine the error in those executions. Hence, PSCLβ and redaction classes are - in some sense - duals of each other. This duality is further expounded apon by Lemma 2.
Lemma 2.
For any given redaction class and any given PSCLβ either:
$\mathbb{1.}$ Every CPTP list in the PSCLβ fits every circuit in the redaction class
$\mathbb{or}$
$\mathbb{2.}$ No CPTP list in the PSCLβ fits any circuit in the redaction class.
I refer to the first of these possibilities as the redaction class and the PSCLβ matching, but also allow the term to apply if the redaction class is replaced by any subset of itself.
Proof.
As in Note 1, if a given CPTP list fits one circuit in a redaction class, it fits every circuit in that redaction class.
As mentioned in Def. 7 if one CPTP list in a given PSCLβ fits a specific circuit then the entire PSCLβ fits that circuit.
The lemma follows from combining these two facts. ∎
II.3 Problem Setting Presentation
I can now present the protocol by which sets of circuits, in my problem setting, are executed, in Protocol 1.
  1. 1.
    Alice provides a set of circuits to execute, $\mathcal{S}$, all within the same redaction class.
  2. 2.
    Bob receives the full details of $\mathcal{S}$, with all single-qubit gates in every element of $\mathcal{S}$ redacted.
  3. 3.
    Bob chooses a SPSCLβ that matches $\mathcal{S}$, denoted $\Xi$, where $\beta\in\mathbb{R}^{+}$ is known to both Alice and Bob.
  4. 4.
    For each circuit, $\mathcal{C}$, in $\mathcal{S}$:
    1. (a)
      Bob chooses a CPTP list, $\xi$, from $\Xi$.
    2. (b)
      $\mathcal{C}$ is executed with $\xi$ determining the error occurring during the execution.
    3. (c)
      Alice and Bob are both given the measurement outcomes of the execution of $\mathcal{C}$.
Protocol 1 Formal Adversarial Model: How Sets of Circuits are Executed
Alice is free to initiate Protocol 1 whenever and as often as she likes; Bob is free to act how he likes to achieve his stated malicious aims but must execute Protocol 1 as prescribed - but is free to act nefariously within the permitted bounds - when it is invoked. While the bulk of this paper focuses on constructing the accreditation protocol protocol, without much regard for Alice, it culminates in Theorem 2 showing that Alice’s problem is solved by the protocol presented herein.
II.4 Physical Origins of the Error Model / Problem Setting
The two main limitations on the adversarial error / Bob requiring justification are:
  1. 1.
    The single-qubit gates are redacted when the circuits are shown to Bob (in step 2 of Protocol 1).
  2. 2.
    Within a single use of Protocol 1, Bob is restricted to only causing error describable by CPTP lists with a single SPSCLβ, for a known $\beta\in\mathbb{R}^{+}$.
The first limitation corresponds to single-qubit gates experiencing gate-independent error. This is a standard assumption the the pre-existing accreditation protocols (in Refs. [22, 23, 24]) and follows from single-qubit gates typically being the least error-prone components of a quantum computer [28, 29] (which has held true over time [30, 31], and can be most clearly seen in Ref [32, Fig.5]): the error is so small that the error in different single-qubit gates does not differ much. This is a “standard [assumption] in the literature on noise characterisation and mitigation” [33] and has seen extensive use in theoretical work [3, 34, 35, 36, 37, 38, 39, 40, 41, 42, 33].
The second limitation is a weakening of the IID assumption in previous accreditation protocols. The intuition behind it is that the same hardware executing very similar circuits in quick succession - as happens in Protocol 1 will experience similar error in each circuit as:
  1. 1.
    The external factors affecting error have to be fluctuating over a very specific time-span - slow enough to not fluctuate much within the execution of circuits but fast enough to change between executions of a circuit - to cause significant differences in the error in different executions executed in quick succession.
  2. 2.
    The hardware executing the similar circuits is the same in each execution and any error-inducing aspects of the hardware are unlikely to change much in the short time between executions.
  3. 3.
    The effect of any aspect of the circuits being executed that may change the error is minimized, as the circuits are very similar.
Although no paper has, to my knowledge, sought to directly validate this limitation, it can also be justified experimentally:
  1. 1.
    Ref. [23, Figure. 5] ran the accreditation protocol - developed therein - many times, producing a probability of error in the traps for each use of the protocol. The set of error probabilities generated by this can be seen to vary little across many uses of the protocol.
  2. 2.
    Ref. [43, Fig. 6, Fig. 7, and Fig. 8] investigated the error rates in single-qubit and two-qubit gates. It found these error rates are very rarely far from their average. Note that the variation shown in this paper is over a much longer time-span (days) than this limitation requires (seconds).
  3. 3.
    Ref. [32, Fig. 9] examined NISQ devices and - across many devices - plotted the error rates of different qubits in each device. Fig. 9 shows the error bars on the error rate to be - with some notable exceptions - small fractions of the error rate and varying much more across qubits than for a fixed qubit over time.
  4. 4.
    Ref. [44, Fig. 1(b)] looked at the error rates of two-qubit gates on differing pairs of qubits, it shows that - again, over a time-span of days - for a specific pair of qubits, the days when the error rate deviates far from its average are rare, with more extreme deviations being rarer.
Finally, this limitation can be seen as an aim of quantum computer hardware engineering: as quantum computers improve and their actual outputs approach the errorless outputs, with decreasing variance, the typical difference - by any measure - between the error contained in two executions of similar circuits will tend to zero.
To my knowledge, I am the first to explicitly state and use this second limitation. However, I note that it is already implicitly accepted in the community by the acceptance of randomized benchmarking [3, 4, 5, 6] as a meaningful measure: if even the same circuit repeated multiple times, in very quick succession, produces wildly varying probabilities of error then the variation of error probabilities implies that previous measures mean almost nothing for future computations or the quality of a quantum device over a meaningful time-span. This is not the case and Ref. [44, Fig. 6] shows that, over a span of days, the results of randomized benchmarking do not vary over a very large range.
There is, however, another, subtle, restriction on Bob: he controls the CPTP maps acting on all executions, but does not control any stochastic elements of the implementation of the CPTP map. I.e. he may specify that a specific CPTP map in the execution applies stochastic Pauli error, such as a dephasing channel, but he does not decide if this results in a Pauli $\hat{Z}$ gate being applied. Furthermore, Bob is not informed of the outcomes of these stochastic processes - in spite of choosing which stochastic processes occur. This represents that the environment - where any memory of past error that the noise represented by Bob has must be - is not, consistently or over long time-spans, protected from noise acting on it and destroying any information/memory it contains, so any hostile intelligence that lurks in the noise must have a comparatively short and small memory  [^8] Quantum hardware engineers have expended much effort to keep the the information inside quantum computers intact, so there is no reason to assume the environment can - with non-negligible probability - store significant amounts of information, for relevant timespans, without any engineering effort to enable it to do so..
III Accreditation Protocol
Sec. III is dedicated to resolving the problem Alice faces in the problem setting established in Sec. II.3. This is equivalent to upgrading an accreditation protocol that assumed error is CPTP and IID (and that single-qubit gates suffer only gate-independent error) to one that works with the error assumptions in Sec. II.3.
III.1 Trap and Target Circuits
In this paper, I do not propose to develop new trap circuits or target circuits. In fact, I would prefer to make minimal changes to the trap and target circuits in Ref. [23]. I will also not regurgitate the exact designs of the trap and target circuits in Ref. [23] and instead - throughout this paper - will assume that I have two efficient classical algorithms, $P_{\textit{targ}}$ and $P_{\textit{trap}}$, that, if given any quantum circuit as input, return the corresponding target and a random [^9] I.e. $P_{\textit{trap}}$ will choose the random gates for the Pauli twirl and error detection and so will return a slightly different circuit each time. matching trap, respectively, of the protocol in Ref. [23].
I briefly note the important features of these trap and target circuits:
  1. 1.
    In target circuits and trap circuits all error occurring is twirled (via Pauli twirls or some other twirl) to stochastic Pauli error.
  2. 2.
    Each execution of a trap acts as a Bernoulli trial, with the success probability - for that trial - being the probability of error occurring in that trap execution.
  3. 3.
    If no error occurs in a trap, it gives a specific output, $m$, and if error does occur the trap does not give the output $m$ with probability at least $k\in[0,1]$.
  4. 4.
    Target circuits and trap circuits differ only in their single-qubit gates.
The above assumptions have mentioned twirling CPTP error to stochastic Pauli error, which I define formally in Def. 8. This is an important step in all accreditation protocols, as it reduces the error to a known, more easily quantified form.
Definition 8.
CPTP error within a quantum circuit is said to be twirled to stochastic Pauli error [45, 46, 47] if, via the addition of only single-qubit Pauli gates to the circuit, the error is effectively transformed to stochastic Pauli error, without otherwise affecting the outputs of the circuit. Likewise, gates are said to be twirled if any error occurring in them is twirled.
As herein noise is now considered to be adversarial, the accreditation protocol of this paper does require a slight modification of the trap circuits and target circuits: Ref. [23] did not consider “hiding” the measurement outcomes of a circuit as there was no adversary to “hide” them from (according to its error model). With adversarial noise (i.e. the problem setting of Sec. II.3), this becomes necessary as otherwise Bob may be able to identify traps and targets based on these outcomes, or base future error on the measurement outcomes of previous circuits. The required hiding / encrypting of measurement outcomes is achieved via Algorithm 2.
$\mathbb{Input:}$
$\bullet$ A circuit, $\mathcal{C}$, to generate traps or a target for.
$\bullet$ Two algorithms, $P_{\textit{trap}}$ and $P_{\textit{targ}}$, as detailed above.
$\bullet$ A Boolean, labelled isTarget, denoting if a trap or target is to be generated.
  1. 1.
    If isTarget == True:
    1. (a)
      $\mathcal{C}^{\prime}$ = $P_{\textit{targ}}\big{(}\mathcal{C}\big{)}$.
    Else:
    1. (a)
      $\mathcal{C}^{\prime}$ = $P_{\textit{trap}}\big{(}\mathcal{C}\big{)}$.
  2. 2.
    Generate a random bit string with the same length as the number of measurements in $\mathcal{C}^{\prime}$, referred to as the key.
  3. 3.
    For measurement, M, in $\mathcal{C}^{\prime}$:
    1. (a)
      Calculate the single-qubit unitary, $\mathcal{U}_{M}$, that, if applied immediately before measurement, M, would flip the outcome.
    2. (b)
      If ( the bit in the key corresponding to measurement M ) == 1:
      1. i.
        Add $\mathcal{U}_{M}$ to $\mathcal{C}^{\prime}$ immediately before measurement M.
$\mathbb{Return}:$ $\mathcal{C}^{\prime}$ and the key.
Algorithm 2 Generating Traps and Targets with Hidden Outputs
I note that the required (by Algorithm 2) single-qubit unitaries, $\mathcal{U}_{M}$, will exist for any single-qubit measurement and that the outcome of the circuit - if the single-qubit gates of the circuit are not known, as is the case for Bob as they are redacted for him - is irretrievable from the measurement outcomes without the key. This hiding of the circuit outputs is information-theoretically secure (i.e. has perfect security) [48], assuming the single-qubit gates are redacted, and comparable to the one-time pads in universal blind quantum computing [49]. Henceforth, I will denote $P_{\textit{targ}}$ and $P_{\textit{targ}}$, with the changes presented above to make their outputs hidden, by $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$ respectively.
III.2 Presentation of the Upgraded Accreditation Protocol
With trap and target circuits established, I can consider the full accreditation protocol to be developed. In line with Sec. III.1, for Sec. III.2, traps and targets will be treated as blackboxes and I will only refer to their construction as being performed by $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$. Their only relevant properties will be that traps detect any error with probability $k\in[0,1]$, all error in a trap or target simulation is reduced to effective Pauli error, and Bob cannot tell the difference between trap and target simulations due to them only differing in their single-qubit gates.
However, before the full protocol can be presented, in Protocol 3, I must establish the statistical foundations of the new protocol - in Sec. III.2.1 - and the core mechanics of the protocol - in Sec. III.2.2. These statistical methods will be used to evaulate the $P_{0}$ that defines Bob’s particular choice of a SPSCLβ using multiple trap executions (still all within a single use of Protocol 1).
III.2.1 Statistical Basis of the New Accreditation Protocol
Before presenting the accreditation protocol, it is first useful to present a purely statistical lemma (Lemma 4) that will later enable the accreditation protocol. This presentation first begins with Def. 9 and Lemma 3.
Definition 9.
For any set of real values , =$\mathcal{R}=\big{\{}r_{j}\big{\}}^{|\mathcal{R}|}_{j=1}$, define: $\underline{\mathrm{Avg}\big{(}\mathcal{R}\big{)}}=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}r_{j}\bigg{)}.$
Lemma 3.
Given a set of real values, $\mathcal{R}\subset[(1-\beta)P_{0},(1+\beta)P_{0}]$; if $\beta\in\mathbb{R}$ is known, $P_{0}$ can be bounded as:
$\dfrac{\mathrm{Avg}(\mathcal{R})}{1+\beta}$ $\leq P_{0}\leq\dfrac{\mathrm{Avg}(\mathcal{R})}{1-\beta}.$ (5)
Proof.
Denote =$\mathcal{R}=\{r_{j}|1\leq j\leq|\mathcal{R}|\}$, then:
$\mathrm{Avg}(\mathcal{R})$ $=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}r_{j}\bigg{)}.$ (6)
As $\forall r_{j}\in\mathcal{R}$, $r_{j}\in[(1-\beta)P_{0},(1+\beta)P_{0}]$; $\forall r_{j}\in\mathcal{R}$, $(1-\beta)P_{0}\leq r_{j}\leq(1+\beta)P_{0}$. Therefore,
$(1-\beta)P_{0}$ $=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}(1-\beta)P_{0}\bigg{)}$ (7)
$\leq\mathrm{Avg}(\mathcal{R})$ (8)
$\leq\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}(1+\beta)P_{0}\bigg{)}=(1+\beta)P_{0}$ (9)
$\Rightarrow\dfrac{\mathrm{Avg}(\mathcal{R})}{1+\beta}$ $\leq P_{0}\leq\dfrac{\mathrm{Avg}(\mathcal{R})}{1-\beta}.$ (10)
I may then present and prove Lemma 4, which makes use of Lemma 3. Whereas Lemma 3 was concerned with characterizing the interval $\mathcal{R}$, Lemma 4 aims to - using Lemma 3 - place a bound on any value occurring within $\mathcal{R}$ by estimating a bound on $P_{0}$.
Lemma 4.
Given a set of real values, $\mathcal{R}\subset[(1-\beta)P_{0},(1+\beta)P_{0}]$; if $\beta\in[0,1)$ is known, $\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$,
$y\leq\bigg{(}1+2\beta\bigg{)}\mathrm{Avg}(\mathcal{R}).$ (11)
Proof.
Using Lemma 3,
$P_{0}\leq\dfrac{\mathrm{Avg}(\mathcal{R})}{1-\beta}.$ (12)
$\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$, $y\leq(1+\beta)P_{0}$. Therefore, $\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$,
$y\leq(1+\beta)P_{0}\leq\dfrac{1+\beta}{1-\beta}\mathrm{Avg}(\mathcal{R}).$ (13)
Using the Taylor series of $\dfrac{1}{1-\beta}$, which holds for all values $\beta$ can take, and neglecting quadratic (or higher) terms:
$\dfrac{1+\beta}{1-\beta}\approx 1+2\beta\Rightarrow y\leq\bigg{(}1+2\beta\bigg{)}\mathrm{Avg}(\mathcal{R}),$ (14)
where I have assumed that the approximation is tight enough to not affect the inequality. ∎
III.2.2 Core Mechanics of the New Accreditation Protocol
Before its formal presentation, in Sec. III.2.3, the core mechanics of the accreditation protocol are presented. This takes the form Lemma 5.
Lemma 5.
If given a classical polynomial algorithm, $P_{\textit{trap}}^{\prime}$, generating traps corresponding to a fixed circuit, $\mathcal{C}$, and the corresponding $m$, and $k$ (as defined in Sec. III.1); assuming the probability of error in each execution is independent of the outcomes of all preceding executions, that any trap detects any specific error with probability at least $k\in[0,1]$, and all traps are executed via Protocol 1, the probability of error in any execution of any circuit executed within the same singular use of Protocol 1 - such as a target circuit  [^10] As the probability of error in the target would, due to the assumptions made, be within $[(1-\beta)P_{0},(1+\beta)P_{0}]$. - as the traps can be upper bounded by:
$\dfrac{1+2\beta}{k}\big{(}\bar{v}+\theta\big{)},$ (15)
using $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps, where: $\bar{v}$ is the fraction of traps giving an incorrect measurement outcome, $\theta\in\mathbb{R}^{+}$ may be chosen arbitrarily, $\beta$ is as in Protocol 1, and $\alpha$ is the confidence required in the bound.
Proof.
Let $\mathcal{R}$ be an ordered set where the $j$th element is the probability of error occurring in the $j$th execution of a set of trap executions interspersed among any number of other circuits. In a singular use of Protocol 1. This implies that the probability of error in any specific trap mentioned above - and in any circuit executed within the same single use of Protocol 1 - is within the interval:
$[P_{0}(1-\beta),P_{0}(1+\beta)].$ (16)
Therefore, due to Lemma 4, for any circuit executed in the aforementioned single use of Protocol 1, labelled circuit $j$, the probability of error occurring in its execution, $p_{j}$, is bounded by:
$p_{j}\leq\bigg{(}1+2\beta\bigg{)}\mathrm{Avg}(\mathcal{R}).$ (17)
The probability a specific trap returns an incorrect measurement outcome, given error occurs, is, by assumption of having the required traps, lower bounded by $k\in[0,1]$.
Therefore, by also using the independence between circuit executions of whether error occurs, $\mathrm{Avg}(\mathcal{R})$ can be approximated - to within error, $\theta$, with confidence $\alpha$ using Hoeffding’s inequality [50] and $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ trap executions.
This approximation of $\mathrm{Avg}(\mathcal{R})$ is simply the fraction of traps returning an incorrect measurement outcome divided by $k$ (to account for the cases where error is not detected). Let $\bar{\nu}$ denote this approximation. Then, with confidence, $\alpha$,
$p_{j}\leq\dfrac{1+2\beta}{k}\big{(}\bar{v}+\theta\big{)}.$ (18)
III.2.3 Formal Presentation of the Accreditation Protocol and Proof of its Correctness
The core components of my accreditation protocol have now been constructed and presented - in Sec. III.2.1 and Sec. III.2.2 - so I can now present my accreditation protocol formally, in Algorithm 3. This algorithm is presented as would be used by Alice, presupposing the problem setting in Sec. II.4.
$\mathbb{Input:}$
$\bullet$ A circuit, $\mathcal{C}$.
$\bullet$ A required accuracy of the bound to output, $\theta$.
$\bullet$ A required confidence in the above accuracy, $\alpha$.
$\bullet$ The relevant $P^{\prime}_{\textit{targ}}$, $P^{\prime}_{\textit{trap}}$, $m$, and $k$.
$\bullet$ The value of $\beta$.
  1. 1.
    Calculate $N_{l}=\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}+1$.
  2. 2.
    $\{\mathcal{C}_{j}\}_{j=1}^{N_{l}}$ = $N_{l}$ circuits generated using $P_{\textit{trap}}$ on input $\mathcal{C}$.
  3. 3.
    $\mathcal{C}^{\prime}$ = circuit generated using $P_{\textit{targ}}$ on input $\mathcal{C}$.
  4. 4.
    Execute every circuit in $\{\mathcal{C}_{j}\}_{j=1}^{N_{l}}\cup\{\mathcal{C}^{\prime}\}$ in a random order using Protocol 1.
  5. 5.
    $\mathrm{TargResult}$ = output of executing $\mathcal{C}^{\prime}$
  6. 6.
    $\mathrm{TrapResult}$= fraction of $\{\tilde{\mathcal{C}}_{j}\}_{j=1}^{N_{l}}$ that does not output $m$.
$\mathbb{Return}:$ $\mathrm{TargResult}$, $\frac{1+2\beta}{k}\big{(}\mathrm{TrapResult}+\theta\big{)}$.
Protocol 3 Formal Accreditation Protocol
The correct functioning of Algorithm 3, as an accreditation protocol, is proven in Theorem 1.
Theorem 1.
Assuming all sets of circuits executed are executed via Protocol 1, and that the classical algorithms, $P^{\prime}_{\textit{targ}}$ and $P^{\prime}_{\textit{trap}}$, generate trap and target circuits with the features listed in Sec. III.1 (i.e. as in Ref. [23]), then Algorithm 3 provides a sample from a probability distribution and a bound on the ideal-actual variation distance between that distribution and the distribution generated by the measurements in $\mathcal{C}$ when no error occurs (to within accuracy $\theta$, with confidence $\alpha$).
Proof.
All circuits executed in Algorithm 3 are executed within a single use of Protocol 1. Therefore, if each probability of error in a given trap is independent, then Lemma 5 implies that, as the target circuit is a circuit executed alongside $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps in a single use of Protocol 1, the probability of error in the target circuit’s execution, $p_{t}$, is upper bounded as:
$p_{t}\leq\bigg{(}\dfrac{1+2\beta}{k}\bigg{)}\big{(}\bar{v}+\theta\big{)},$ (19)
where $\bar{v}$ is the fraction of traps returning an incorrect measurement result and $k\in[0,1]$ is as defined in Sec. III.1. As mentioned, for Eqn. 19 to hold, it is required that the probability of error in each trap is independent of the outcomes of every prior trap. This is not initially guaranteed as Bob is free to choose probabilities based on previous measurement outcomes. However, due to the changes made to each trap and target in Algorithm 2, to hide the outputs of each trap from Bob with perfect security, Bob does not know the result of any trap or target. Therefore, he cannot make any of his choices based on the results of any prior trap or target  [^11] I additionally note that Bob cannot remove this independence by basing his choice of CPTP maps on error in previous traps. Although Bob knows which CPTP maps are applied to all previous circuits executed, he does not know which errors have occurred due to the stochastic processes involved in the CPTP maps he chose, only the probabilities of each occurring as he doesn’t control any aspect of the CPTP maps’ applications, due to the construction of Protocol 1. Therefore, Bob cannot base the choice of CPTP maps in a trap on the outcome of any stochastic property of previous traps as they are unknown to him.. Hence, each trap can be considered to be independent i.e. the probability of error in each trap or target is the result of Bob’s choices but must remain completely independent of the results of any prior trap or target.
As per Protocol 1, when Bob is given the set of circuits to execute - which in this case would be the traps with the target hidden among them - the single-qubit gates are redacted: therefore, as the targets and the traps only differ in their single-qubit gates (as assumed in Sec. III.1) Bob cannot distinguish traps from targets at any stage in Protocol 3. The redaction additionally allows for the unbreakable encryption of the output of each trap or target, as in Algorithm 2, so Bob cannot locate the target via the outputs either. ∎
Theorem 1 is a proof of the correctness of the protocol developed herein, but the main aim of this paper is to provide Alice with a solution to her predicament outlined in Sec. II.1 and Sec. II.3. To this end, I present another theorem - without proof - that follows directly from Theorem 1 catering to Alice more directly.
Theorem 2.
Protocol 3 allows Alice to get Bob to implement any quantum circuit of her choosing and be able to place a bound on the ideal-actual variation distance (with arbitrary confidence $\alpha$) of the results she receives.
IV Discussion
In this paper I have upgraded the accreditation protocol of Ref. [23] to consider all error to be adversarial. This has necessitated the development of a new model of adversarial error, where Bob is limited according to experimentally extracted rules. These rules - that single-qubit gates are hidden from Bob and that the probability of error in similar circuits executed in quick succession have similar probabilities of error - allow my upgraded accreditation protocol to function almost identically when experiencing adversarial error as the one in Ref. [22] does when the error is IID. Therefore, the desirable qualities the protocol from Ref. [22], such as not requiring lengthening the circuit to accredit, or adding excessively many extra single-qubit gates, or any extra two-qubit gates, or ancilla qubits, or trusting any aspect of a computation, or extra connectivity between qubits are preserved; leading to no diminution in its suitability for near-term usage.
The principle question left open is to what extent can the limits on Bob be relaxed. Relaxing the limitations would have to entail changes to Protocol 1. I believe the redaction of single-qubit gates cannot be eliminated without substantial changes to the trap, as otherwise Bob could distinguish target and trap circuits - although it is not clear trap and target circuits that are indistinguishable to Bob can be constructed without requiring another limitation on Bob.
A slight improvement can be obtained using Ref. [23, Theorem 1], which allows the error in single-qubit gates to be weakly gate dependent: this could be added to my cryptographic model by allowing Bob to define error in the - still redacted - single-qubit gates that depends on the single-qubit gates  [^12] Single-qubit gates would still have to be redacted to prevent Bob identifying the target but he could - perhaps - specify some dependence without seeing the single-qubit gates e.g. “if that gate is an $\hat{X}$ gate…” but the differing CPTP maps - for each single-qubit gate - must differ, in the diamond norm, by at most some small known value. This is a further step towards reality and the restriction on the differences between the gate-dependent error reflects the very small errors in single-qubit gates [32, Fig.5].
Finally I note that, if I take the philosophical position that error that does not affect the state of the qubits, or only adds an overall phase, effectively does not occur and can be considered to have never been applied, I can relax the assumption that Bob does not directly control stochastic processes in the CPTP maps modelling error. This is as the Pauli twirling, accompanied by the random Hadamard gates, make it impossible for Bob to know whether gates he has applied have any affect - which, as mentioned above, can be interpreted as him not being able to decide if he has effectively applied the error. This breaks any ability of Bob to apply error in a way that is dependent on the error in previous circuit executions. Due to this, the error can be considered to be independent between circuit executions.
V Acknowledgements
I would like to thank Animesh Datta and Theodoros Kapourniotis for useful conversations.
This work was supported, in part, by a EPSRC IAA grant (G.PXAD.0702.EXP) and the UKRI ExCALIBUR project QEVEC (EP/W00772X/2).